Twice The sum of 2 consecutive odd numbers is 22 more than their difference. What are the numbers?
Answer by TimothyLamb(4379)

---
2(x + y) = y - x + 22
y = x + 2
---
2x + 2y = y - x + 22
x - y = -2
---
3x + y = 22
x - y = -2
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 5
y = 7
---
system matrix
3 1
1 -1
---
inverse of system matrix
0.25 0.25
0.25 -0.75
---
determinant of system matrix = -4
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php